# Corks, exotic 4-manifolds and genus functions

### v3

@@ 10. Exotically knotted submanifolds
Let $n$ be a non-negative even integer, and set $r=2/n$. In the $n=0$ case, the claim immediately follows from Corollary 10.9, so we prove the $n>0$ case. Let $K$ be an invertible knot in $S^{3}$ with $|\sigma(K)|\geq 2n-1$, where $\sigma(K)$ denotes the signature of $K$. Let $Y_{1}$ be the 4-manifold obtained from $D^{4}$ by attaching a 2-handle along $K$ with 00-framing, and let $M_{r}$ be the boundary of the $r$-fold boundary sum $Y_{r}$ of $Y_{1}$. Note that $M_{r}$ is the $r$-fold connected sum of $M_{1}$. This 3-manifold $M_{r}$ admits an embedding into the closed 4-manifold $r(S^{2}\times S^{2})$ by taking the double of $Y_{r}$. By Theorem 10.7, we obtain a closed 3-manifold $M^{\prime}_{r}$ homology cobordant to $M_{r}$ such that $r(S^{2}\times S^{2})$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M^{\prime}_{r}$. Note that $r(S^{2}\times S^{2})$ is simply connected and satisfies $b_{2}=n$.
We will show that $M^{\prime}_{r}$ does not admit any embedding into a closed 4-manifold with $b_{2}<n$ as a separating submanifold. Since $M^{\prime}_{r}$ is homology cobordant to $M_{r}$, we have a homology cobordism from $M^{\prime}_{r}$ to $M^{\prime}_{r}$ that contains $M_{r}$ in the interior as a separating submanifold. Now suppose to the contrary that $M^{\prime}_{r}$ is diffeomorphic to a separating submanifold of a closed 4-manifold with $b_{2}<n$. We remove the tubular neighborhood of $M^{\prime}_{r}$ from the closed 4-manifold and glue the homology cobordism from $M^{\prime}_{r}$ to $M^{\prime}_{r}$ along the resulting boundary. As seen from the Mayer–Vietoris exact sequence, the resulting closed 4-manifold satisfies $b_{2}<n$. Clearly, $M_{r}$ is embedded in the resulting closed 4-manifold as a separating submanifold. This contradicts a result of Kawauchi that $M_{r}$ admits no such an embedding (see [43, the proof of Theorem 2.5]). Therefore, the desired claim follows. ∎
@@ 11. Diffeomorphism invariants of genus function type
In this section, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds and give examples of such invariants which can detect exotic 4-manifolds. Furthermore, we show that an invariant of genus function type yields lower bounds for values of genus functions and shares many properties of genus functions. Throughout this section, we allow boundaries of $4$-manifolds to be disconnected.
@@ 11.1. Definitions and examples
To state the notion, we consider the category of 4-manifolds whose morphisms are orientation preserving diffeomorphisms. Note that manifolds are oriented due to our convention. For a fixed ordered set $R$, we also consider the category of $R$-valued maps whose morphisms are maps between the source sets of $R$-valued maps.
@@
Let $R$ be an ordered set. A functor $G$ from the category of 4-manifolds to the category of $R$-valued maps will be called an oriented diffeomorphism invariant of genus function type if the following conditions hold.
@@
    $G$ assigns to each $4$-manifold $X$ a map $G_{X}:H_{2}(X)\to R$, which will be called the $G$ invariant of $X$.
@@
    Let $X,Y$ be arbitrary 4-manifolds. For any orientation preserving embedding $\varphi:X\to Y$ and any class $\alpha\in H_{2}(X)$, the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ satisfies $G_{X}(\alpha)\geq G_{Y}(\varphi_{*}(\alpha))$.
If furthermore $G$ satisfies the condition (ii) for any embedding $\varphi$ that does not necessarily preserves the orientations, then $G$ will be called a ((((unoriented)))) diffeomorphism invariant of genus function type.
@@
Let $G$ be a (possibly oriented) diffeomorphism invariant of genus function type, and let $X,Y$ be $4$-manifolds.
(1) We will say that a continuous map $\varphi:X\to Y$ preserves $G$ invariants if the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the values of the $G$ invariants.
(2) The $G$ invariants of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $\varphi:X\to Y$ that preserves the $G$ invariants. If $\varphi$ furthermore preserves the orientations, then the $G$ invariants of $X$ and $Y$ will be called orientedly equivalent.
(3) The $G$ invariants of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the intersection forms and $G$ invariants.
Due to Remark 11.2, $G$ invariants of orientedly diffeomorphic 4-manifolds are orientedly equivalent. Also, if $G$ invariants of two 4-manifolds are orientedly equivalent, then they are algebraically equivalent. Furthermore, if $G$ is an unoriented diffeomorphism invariant of genus function type, then $G$ invariants of unorientedly diffeomorphic 4-manifolds are equivalent.
We also introduce the following terminologies, extracting properties of genus functions.
@@
(1) Let $R$ be an additive semigroup and suppose that the $G$ invariant of each 4-manifold is an $R$-valued map. We will say that $G$ is subadditive with respect to gluing if $G$ satisfies the following property for any 4-manifold $X$ with non-empty (possibly disconnected) boundary $M$ and for any cobordism $P$ from $M$ to a (possibly disconnected) closed 3-manifold. Set $X^{\prime}=X\cup_{M}P$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Then the inequality $G_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq G_{X}(\alpha)+G_{P}(\beta)$ holds for any $\alpha\in H_{2}(X)$ and any $\beta\in H_{2}(P)$.
(2) We will say that $G$ is preserved under attachment of a $3$-handle ((((resp. a $4$-handle)))) if the following property holds. For any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching a 3-handle (resp. a 4-handle), the inclusion $X\hookrightarrow X^{\prime}$ preserves the $G$ invariants.
Genus functions clearly have the above subadditivity and are preserved by 3- and 4-handles due to Lemma 7.2. We remark that all the properties of genus functions used for the proofs of the results in Sections 7 and 8 are only those in Definitions 11.1 and 11.4. We also define “torsion-free $G$ invariants” and related terminologies in the same way as the torsion-free genus functions.
The genus functions have a property that $g_{X}(0)=g_{D^{4}}(0)$ for any 4-manifold $X$. This property gives the following simple obstructions for having the properties on 3- and 4-handles. As we will see in Example 11.6, these obstructions are effective.
Lemma 11.5.
$(1)$ If $G$ is preserved under attachment of a $4$-handle, then $G_{D^{4}}(0)=G_{S^{4}}(0)$.
$(2)$ If $G$ is preserved under attachment of a $3$-handle, then every $\alpha\in H_{2}(S^{2}\times D^{2})$ satisfies $G_{S^{2}\times D^{2}}(\alpha)=G_{D^{4}}(0)=G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)$.
Proof.
The claim (1) and the equality $G_{S^{2}\times D^{2}}(\alpha)=G_{D^{4}}(0)$ are straightforward. Since attachment of a 1-handle preserves $G$ invariants due to Remark 11.11, the equality $G_{D^{4}}(0)=G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)$ follows. ∎
Other than genus functions, we give the following examples of genus function type which can detect exotic 4-manifolds.
@@
(1) For a 4-manifold $X$, we define the map $G_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$ as follows. Suppose $X$ admits an orientation preserving embedding $\varphi:X\to Z$ into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ having a non-vanishing Seiberg–Witten invariant. We furthermore assume $Z$ is of simple type in the sense of [56]. Let $K\in H^{2}(Z)$ be a Seiberg-Witten basic class of $Z$. Then, for $\alpha\in H_{2}(X)$, we define $G_{X}(\alpha)$ as the maximal value of the integer $|\langle K,\varphi_{*}(\alpha)\rangle|$ for all $Z$, $\varphi$ and $K$. Note that $G_{X}(\alpha)<\infty$ for any $\alpha$ due to the adjunction inequality ([46], [56]). In the case where $X$ admits no embedding into such a closed 4-manifold $Z$, we set $G_{X}(\alpha)=-\infty$ for any $\alpha\in H_{2}(X)$. It is easy to see that $G_{X}$’s form an oriented diffeomorphism invariant $G$ of genus function type. This invariant can detect various families of orientedly exotic 4-manifolds, as implicitly shown in [11], [12], [67], [68] and [70]. We can define several variations of the above invariant $G$ by using the Seiberg–Witten invariant and the Bauer–Furuta invariant. These invariants will be further discussed in [72].
The above $G$ clearly has the subadditivity with respect to gluing, but it is preserved under an attachment of neither 3- nor 4-handles. The claim for a 4-handle follows from the equalities $G_{D^{4}}(0)=0$ and $G_{S^{4}}(0)=-\infty$ due to Lemma 11.5. Similarly, the claim for a 3-handle follows from the equalities $G_{D^{4}}(0)=0$ and $G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)=-\infty$ due to Lemma 11.5. Note that for any closed 4-manifold $Z$ with $b_{2}^{+}>1$, the Seiberg–Witten invariant of $Z\#(S^{1}\times S^{3})$ vanishes, since the parities of $b_{2}^{+}-b_{1}$ of these two 4-manifolds are distinct.
@@
(2) Recently, Ren and Willis [58] introduced an interesting new invariant of 4-manifolds called the lasagna $s$-invariant by using skein lasagna modules. The lasagna $s$-invariant (with empty boundary link) of a 4-manifold $X$ is a map $s_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$. According to their result [58, Proposition 1.11], this invariant is an oriented diffeomorphism invariant of genus function type. Furthermore, they showed that this invariant can detect many orientedly exotic pairs of 4-manifolds.
@@
(3) Let $G$ be an oriented diffeomorphism invariant of genus function type. We can modify $G$ into an unoriented invariant, for example, as follows. For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $G^{\prime}_{X}(\alpha)=\max\{G_{X}(\alpha),G_{\overline{X}}(\alpha)\}$. The resulting $G^{\prime}$ is an unoriented diffeomorphism invariant of genus function type. For example, in the case of the invariant defined in (1), the unoriented version can distinguish various families of unorientedly exotic 4-manifolds as implicitly shown in [11] and [68].
@@ 11.2. Properties
We first show that $G$ yields a function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ that gives lower bounds for values of genus functions. For integers $g,n$ with $g\geq 0$, let $S(g,n)$ be the total space of the $D^{2}$-bundle over the closed surface of genus $g$ with Euler number $n$, and let $\gamma(g,n)$ denote a generator of $H_{2}(S(g,n))(\cong\mathbb{Z})$. For $(r,n)\in R\times\mathbb{Z}$, we define an element $A_{G}(r,n)$ of ${}\mathbb{Z}\cup\{\infty\}$ as follows. In the case where $G_{S(g,n)}(\gamma(g,n))<r<G_{S(g+1,n)}(\gamma(g+1,n))$ for some integer $g\geq 0$, we set $A_{G}(r,n)=g+1$. In the case where $r=G_{S(g,n)}(\gamma(g,n))$ for some integer $g\geq 0$, we set
In the case where $r<G_{S(g,n)}(\gamma(g,n))$ (resp. $r>G_{S(g,n)}(\gamma(g,n))$) for any integer $g\geq 0$, we set $A_{G}(r,n)=0$ (resp. $A_{G}(r,n)=\infty$). The resulting function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ is well-defined due to the following lemma.
@@
As easily seen from a handlebody diagram, the 4-manifold $S(g,n)$ admits an orientation-preserving self-diffeomorphism whose induced isomorphism maps $\gamma(g,n)$ to $-\gamma(g,n)$. The claim (1) thus follows. Due to the assumption $g_{2}\geq g_{1}$, we easily see that $\gamma(g_{1},n)$ is represented by a closed surface of genus $g_{2}$ with self-intersection number $n$, which is obtained from the base surface of $S(g_{1},n)$ by taking a connected sum with a locally embedded surface. We thus have an embedding of $S(g_{2},n)$ into $S(g_{1},n)$ whose induced homomorphism maps $\gamma(g_{2},n)$ to $\gamma(g_{1},n)$ or to $-\gamma(g_{1},n)$. The claim (2) thus follows from the definition of genus function type and the claim (1). ∎
@@
In the above claim, $\alpha\cdot\alpha$ denotes the self-intersection number of $\alpha$. We remark that this inequality can be sharp, depending on the invariant $G$. For example, for the invariant in Example 11.6.(1) defined by using the Seiberg–Witten invariant, there are many examples of 4-manifolds and their homology classes for which the above bound gives sharp estimates, realizing arbitrary large values of genus functions. This claim can be easily checked using the adjunction inequality for the Seiberg–Witten invariant.
@@
We set $g_{0}=g_{X}(\alpha)$. Then $\alpha$ is represented by a closed surface of genus $g_{0}$ with the self-intersection number $n$. We thus have an embedding of $S(g_{0},n)$ into $X$ whose induced homomorphism maps $\gamma(g_{0},n)$ to $\alpha$, showing that $G_{S(g_{0},n)}(\gamma(g_{0},n))\geq G_{X}(\alpha)$. This inequality and the definition of $A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$ shows $g_{0}\geq A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$. The claim thus follows. ∎
We note that for each fixed $n$, the function $A_{G}(\;,n):R\to\mathbb{Z}$ is non-decreasing. We observe that this function is the best possible one among similar functions derived from $G$ in the following sense.
Proposition 11.9.
Let $n$ be an integer. Suppose a non-increasing function $A:R\to\mathbb{Z}\cup\{\infty\}$ satisfies $g_{X}(\alpha)\geq A(G_{X}(\alpha))$ for any $4$-manifold $X$ and any $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$. Then the inequality $A_{G}(G_{X}(\alpha),n)\geq A(G_{X}(\alpha))$ holds for any $4$-manifold $X$ and any $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$.
Proof.
Suppose to the contrary that the inequality $A_{G}(G_{X}(\alpha),n)<A(G_{X}(\alpha))$ holds for some 4-manifold $X$ and some $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$. We set $g_{A}=A(G_{X}(\alpha))$. Then, we have $g_{A}>A_{G}(G_{X}(\alpha),n)$, which implies $G_{S(g_{A}-1,n)}(\gamma(g_{A}-1,n))\geq G_{X}(\alpha)$. By the assumptions on $A$, this inequality implies $g_{A}-1\geq A(G_{S(g_{A}-1,n)}(\gamma(g_{A}-1,n)))\geq g_{A}$, giving a contradiction. Therefore, the claim follows. ∎
@@
Let $M,N$ be ((((possibly disconnected)))) closed $3$-manifolds, let $P$ be a cobordism from $M$ to $N$, and let $X$ be a $4$-manifold with $\partial X=M$. Then the following hold.
$(1)$ If $P$ is invertible, then the inclusion $X\hookrightarrow X\cup_{M}P$ preserves the $G$ invariants.
$(2)$ Suppose that $G$ is preserved by attachments of $3$- and $4$-handles. If $P$ is quasi-invertible, then the inclusion $X\hookrightarrow X\cup_{M}P$ preserves the $G$ invariants.
@@
We can prove these claims similarly to Theorem 7.1. See also Remark 7.3. ∎

### v5

@@ 10. Exotically knotted submanifolds
We set $r=n/2\geq 0$ and $k=m-r\geq 0$. In the $n=0$ case, the claim follows from Corollary 10.9, so we prove the $n>0$ case. Let $K$ be an invertible knot in $S^{3}$ with $|\sigma(K)|\geq 2n-1$, where $\sigma(K)$ denotes the signature of $K$. For example, many torus knots satisfy this condition. Let $Y$ be the 4-manifold obtained from $D^{4}$ by attaching a 2-handle along $K$ with 0-framing, and let $Y_{r,k}$ be the boundary sum of $r$ copies of $Y$ and $k$ copies of $S^{2}\times D^{2}$. We denote by $M_{r,k}$ the boundary $\partial Y_{r,k}$. We can embed each 4-manifold $Y_{r,k}$ and hence $M_{r,k}$ into the closed 4-manifold $r(S^{2}\times S^{2})$ by taking the double of $Y_{r,0}$, which contains $Y_{r,k}$ as a submanifold.
By Theorem 10.7, we obtain a closed 3-manifold $M^{\prime}_{r,k}$ homology cobordant to $M_{r,k}$ such that $r(S^{2}\times S^{2})$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M^{\prime}_{r,k}$. As easily seen, we have $b_{1}(M^{\prime}_{r,k})=b_{1}(M_{r,k})=r+k=m$. Note that $r(S^{2}\times S^{2})$ is simply connected and satisfies $b_{2}=2r=n$.
Since $M^{\prime}_{r,k}$ is homology cobordant to $M_{r,k}$, we have a homology cobordism $P_{r,k}$ from $M^{\prime}_{r,k}$ to $M_{r,k}=M_{r,0}\#k(S^{2}\times S^{1})$. By attaching 3-handles to $P_{r,k}$, we obtain a cobordism from $M^{\prime}_{r,k}$ to $M_{r,0}$. By further attaching 1-handles to this cobordism and then attaching the upside down cobordism of $P_{r,k}$, we obtain a cobordism $P^{\prime}_{r,k}$ from $M^{\prime}_{r,k}$ to itself whose interior contains $M_{r,0}$ as a separating submanifold.
Now we show that each $M^{\prime}_{r,k}$ admits no embedding into a closed 4-manifold with $b_{2}<n$ as a separating submanifold. Suppose to the contrary that $M^{\prime}_{r,k}$ is diffeomorphic to a separating submanifold of a closed 4-manifold with $b_{2}<n$. We remove the tubular neighborhood of the submanifold from the closed 4-manifold and glue the cobordism $P^{\prime}_{r,k}$ along the resulting boundary. The resulting closed 4-manifold still satisfies $b_{2}<n$ as follows from the Mayer–Vietoris exact sequence, and furthermore it contains $M_{r,0}$ as a separating submanifold. This contradicts a result of Kawauchi that $M_{r,0}$ admits no such embedding ([44, the proof of Theorem 2.5]). Therefore, the desired claim follows. ∎
As shown in the above proof, the simply connected closed $4$-manifold with $b_{2}=n$ in Corollary 10.10 can be taken as $\frac{n}{2}(S^{2}\times S^{2})$ for each $n$.
@@ 11. Diffeomorphism invariants of genus function type
In this section, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds and give examples of such invariants which can detect exotic 4-manifolds. Furthermore, we show that any invariant of genus function type yields lower bounds for values of genus functions and shares many properties of genus functions. Throughout this section, we allow boundaries of $4$-manifolds to be disconnected.
@@ 11.1. Definitions and examples
To state the notion, we consider the category of 4-manifolds whose morphisms are orientation-preserving diffeomorphisms. Note that manifolds are oriented due to our convention. For a fixed ordered set $R$, we also consider the category of $R$-valued maps whose morphisms are maps between the source sets of $R$-valued maps.
@@
Let $R$ be an ordered set. A functor $G$ from the category of 4-manifolds to the category of $R$-valued maps is called an oriented diffeomorphism invariant of genus function type if the following conditions hold.
@@
    $G$ assigns to each $4$-manifold $X$ a map $G_{X}:H_{2}(X)\to R$, which we call the $G$ invariant of $X$.
@@
    Let $X,Y$ be arbitrary 4-manifolds. For any orientation-preserving embedding $\varphi:X\to Y$ and any class $\alpha\in H_{2}(X)$, the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ satisfies $G_{X}(\alpha)\geq G_{Y}(\varphi_{*}(\alpha))$.
If furthermore $G$ satisfies the condition (ii) for any embedding $\varphi$ that does not necessarily preserves the orientations, then $G$ is called a ((unoriented)) diffeomorphism invariant of genus function type.
@@
Let $G$ be an oriented diffeomorphism invariant of genus function type, and let $X,Y$ be $4$-manifolds.
(1) We will say that a continuous map $\varphi:X\to Y$ preserves the $G$ invariants if the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the values of the $G$ invariants.
(2) The $G$ invariants of $X$ and $Y$ are called equivalent if there exists a homeomorphism $\varphi:X\to Y$ that preserves the $G$ invariants. If $\varphi$ furthermore preserves the orientations, then the $G$ invariants of $X$ and $Y$ are called orientedly equivalent.
(3) The $G$ invariants of $X$ and $Y$ are called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the intersection forms and $G$ invariants.
Due to Remark 11.2, $G$ invariants of orientedly diffeomorphic 4-manifolds are orientedly equivalent. Also, if $G$ invariants of two 4-manifolds are orientedly equivalent, then they are algebraically equivalent. Furthermore, if $G$ is an unoriented diffeomorphism invariant of genus function type, then $G$ invariants of (unorientedly) diffeomorphic 4-manifolds are equivalent. We also define “torsion-free $G$ invariants” and related terminologies in the same way as for the torsion-free genus functions.
We introduce the following terminologies extracting properties of genus functions.
@@
(1) Let $R$ be an additive magma (i.e., a set equipped with an additve binary operation ++) and suppose that the $G$ invariant of each 4-manifold is an $R$-valued map. We will say that $G$ is subadditive with respect to gluing if $G$ has the following property. Let $X^{\prime}$ be a 4-manifold obtained from a 4-manifold $X$ by attaching a 4-dimensional cobordism $P$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Then every $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(P)$ satisfy $G_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq G_{X}(\alpha)+G_{P}(\beta)$.
(2) We will say that $G$ is preserved under attachment of a $3$-handle ((resp. a 44-handle)) if for any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching a 3-handle (resp. a 4-handle), the inclusion $X\hookrightarrow X^{\prime}$ preserves the $G$ invariants.
(3) Let $P$ be a 4-dimensional cobordism. We will say that $G$ is preserved under attachment of $P$ if for any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching the cobordism $P$, the inclusion $X\hookrightarrow X^{\prime}$ preserves the $G$ invariants.
(4) Let $Z$ be a $4$-manifold. We will say that $G$ is preserved under connected sum with $Z$ if for any 4-manifold $X$, the inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the $G$ invariants of $X$ and X#ZX\#Z. In the case where $\partial Z$ is non-empty, we will say that $G$ is preserved under boundary sum with $Z$ if for any 4-manifold $X$ with non-empty boundary and for any choice of a boundary sum $X\natural Z$, the inclusion $X\hookrightarrow X\natural Z$ preserves the $G$ invariants.
The genus function clearly has the above subadditivity and is preserved under attachment of a 3- and a 4-handle due to Lemma 7.2. All the properties of the genus function used in the proofs of the results in Sections 7 and 8 are only those in Definitions 11.1 and 11.4. It might be worth to remark that the genus function furthermore has the property that every 4-manifold $X$ satisfies $g_{X}(0)=g_{D^{4}}(0)$.
We give other examples of oriented diffeomorphism invariants of genus function type which can detect exotic 4-manifolds.
@@
(1) For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, let $E^{SW}_{X}(\alpha)$ be the maximum value of an integer $|\langle K,\varphi_{*}(\alpha)\rangle|$, where $\varphi:X\to Z$ is an orientation-preserving embedding of $X$ into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ having non-vanishing Seiberg–Witten invariant, and $K$ is a Seiberg–Witten basic class of $Z$. We furthermore assume $Z$ is of Seiberg–Witten simple type in the sense of [58] so that the adjunction inequalities ([48], [58]) guarantee $E^{SW}_{X}(\alpha)<\infty$. In the case where $X$ admits no embedding into such a closed 4-manifold $Z$, we set $E^{SW}_{X}(\alpha)=$-$\infty$. Then, the resulting map $E^{SW}_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$ gives an oriented diffeomorphism invariant of genus function type by varying $X$.
Algebraic equivalent classes of this invariant ${E}^{SW}$ can detect various families of orientedly exotic 4-manifolds, as implicitly shown in [10], [11], [12], [69], [70] and [72]. This invariant ${E}^{SW}$ clearly has the subadditivity with respect to gluing, but it is not preserved under attachment of a 4-handle, since $E^{SW}_{D^{4}}(0)=0$ and $E^{SW}_{S^{4}}(0)=$-$\infty$. We can define several variations of this invariant by using the Seiberg–Witten invariant and the Bauer–Furuta invariant. These invariants will be further discussed in [74].
@@
(2) Recently, Ren and Willis [60] introduced an interesting new invariant of 4-manifolds called the lasagna $s$-invariant by using skein lasagna modules. The lasagna $s$-invariant (with empty boundary link) of a 4-manifold $X$ is a map $s_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$. According to their result [60, Proposition 1.11], this invariant is an oriented diffeomorphism invariant of genus function type. Furthermore, they showed that this invariant can detect many orientedly exotic pairs of 4-manifolds.
(3) In the rest of this section, an immersed 2-sphere means the one whose singular points are only finitely many transverse double points. For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, let $g^{\mathrm{dp}}_{X}(\alpha)$ be the minimum number of double points of an immersed 2-sphere in $X$ representing $\alpha$. In the case where $\alpha$ is not represented by any immersed 2-sphere, we set $g^{\mathrm{dp}}_{X}(\alpha)=\infty$. Then, the resulting map $g^{\mathrm{dp}}_{X}:H_{2}(X)\to\mathbb{Z}\cup\{\infty\}$ gives an unoriented diffeomorphism invariant of genus function type by varying $X$. Similarly to the genus function, determining the values of $g^{\mathrm{dp}}_{X}$ has been an interesting problem, and it is in general quite hard (see e.g [29], [61]).
As is well-known, in the case where $X$ is simply connected, every $\alpha\in H_{2}(X)$ satisfies $g^{\mathrm{dp}}_{X}(\alpha)<\infty$ (see [39, Remark 1.2.4.]). By contrast, in the case where $\pi_{2}(X)=0$, every non-zero class $\alpha\in H_{2}(X)$ satisfies $g^{\mathrm{dp}}_{X}(\alpha)=\infty$. In either case, every $\alpha\in H_{2}(X)$ satisfies $g^{\mathrm{dp}}_{X}(\alpha)\geq g_{X}(\alpha)$, since any double point can be eliminated by increasing the genus (see e.g. [39, p.38]). Using this fact, we easily see that algebraic equivalent classes of $g^{\mathrm{dp}}$ can detect many of infinite families of exotic 4-manifolds in [12], [69], [72]. The invariant $g^{\mathrm{dp}}$ clearly has the subadditivity with respect to gluing, and we can show that this invariant is preserved under attachment of a 3- and a 4-handle, similarly to the proof of Lemma 7.2.
@@
(4) Let $G$ be an oriented diffeomorphism invariant of genus function type. We can modify $G$ into an unoriented invariant, for example, as follows. For a 4-manifold $X$, we set $\overline{G}_{X}=G_{\overline{X}}$. Note that the resulting orientation reversal invariant $\overline{G}$ is also of genus function type. For a class $\alpha\in H_{2}(X)$, we set $|G|_{X}(\alpha)=\max\{G_{X}(\alpha),G_{\overline{X}}(\alpha)\}$. The resulting $|G|$ is an unoriented diffeomorphism invariant of genus function type. For example, in the case of the invariant $E^{SW}$ defined in (1), the unoriented version $|E^{SW}|$ can distinguish various families of unorientedly exotic 4-manifolds as implicitly shown in [11] and [70].
@@ 11.2. Properties
We first show that $G$ yields a function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ that gives lower bounds for values of genus functions. For integers $g,n$ with $g\geq 0$, let $S(g,n)$ be the total space of the $D^{2}$-bundle over the closed surface of genus $g$ with Euler number $n$, and let $\gamma(g,n)$ denote a generator of $H_{2}(S(g,n))(\cong\mathbb{Z})$. To define the function $A_{G}$, we observe the following lemma.
@@
As easily seen from a handlebody diagram, each $S(g,n)$ admits an orientation-preserving self-diffeomorphism whose induced isomorphism maps $\gamma(g,n)$ to $\gamma(g,n)$. The claim (1) thus follows. Due to the assumption $g_{2}\geq g_{1}$, we easily see that $\gamma(g_{1},n)$ is represented by a closed surface of genus $g_{2}$ with self-intersection number $n$, which is obtained from that of genus $g_{1}$ by taking a connected sum with a locally embedded surface. We thus have an embedding of $S(g_{2},n)$ into $S(g_{1},n)$ whose induced homomorphism maps $\gamma(g_{2},n)$ to $\gamma(g_{1},n)$ or to $\gamma(g_{1},n)$. The claim (2) thus follows from the definition of genus function type and the claim (1). ∎
For $(r,n)\in R\times\mathbb{Z}$, we define an element $A_{G}(r,n)$ of $\mathbb{Z}\cup\{\infty\}$ as follows. In the case where $r=G_{S(g,n)}(\gamma(g,n))$ for some integer $g\geq 0$, we set
In the case where $G_{S(g,n)}(\gamma(g,n))<r<G_{S(g+1,n)}(\gamma(g+1,n))$ for some integer $g\geq 0$, we set $A_{G}(r,n)=g+1$. In the case where $r<G_{S(g,n)}(\gamma(g,n))$ (resp. $r>G_{S(g,n)}(\gamma(g,n))$) for any integer $g\geq 0$, we set $A_{G}(r,n)=0$ (resp. $A_{G}(r,n)=\infty$). The resulting function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ is well-defined due to Lemma 11.6.
@@
In the above claim, $\alpha\cdot\alpha$ denotes the self-intersection number of $\alpha$.
@@
We set $g_{0}=g_{X}(\alpha)$. Then $\alpha$ is represented by a closed surface of genus $g_{0}$. Let $n$ be its self-intersection number. We have an embedding of $S(g_{0},n)$ into $X$ whose induced homomorphism maps $\gamma(g_{0},n)$ to $\alpha$, showing that $G_{S(g_{0},n)}(\gamma(g_{0},n))\geq G_{X}(\alpha)$. This inequality and the definition of $A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$ shows $g_{0}\geq A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$. The claim thus follows. ∎
The inequality in Proposition 11.7 can be sharp, depending on the invariant $G$. For example, in the case of the genus function, the equality always holds. For the invariant $E^{SW}$ defined in Example 11.5.(1) by using the Seiberg–Witten invariant, we can easily determine all the values of the function $A_{E^{SW}}$ using the adjunction inequality. Furthermore, we easily see that there are many examples of 4-manifolds and their homology classes for which the above bound is sharp.
@@
$(1)$ $G$ is preserved under attachment of any invertible cobordism and also under boundary sum with any codimension-0 submanifold of $S^{4}$ having non-empty boundary.
$(2)$ If $G$ is preserved under attachment of a $4$-handle, then $G$ is preserved under connected sum with any codimension-0 submanifold of $S^{4}$.
$(3)$ Suppose $G$ is preserved under attachment of a $3$- and a $4$-handle. Then, $G$ is preserved under attachment of any quasi-invertible cobordism. Furthermore, for any $S^{2}$-link $L$, the invariant $G$ is preserved under boundary sum with any codimension-0 submanifold of $S^{4}(L)$ having non-empty boundary and also under connected sum with any codimension-0 submanifold of $S^{4}(L)$.
@@
We can prove these claims similarly to the proofs of Theorems 7.1 and 8.4. See also Remark 7.3. ∎